OpenStudy (anonymous):

Question 7 (Multiple Choice Worth 1 points) (08.01 MC) Lines C and D are represented by the equations given below: Line C: y = x + 8 Line D: y = 3x + 2 Which of these shows the solution to the system of equations and explains why? (3, 11), because the point does not lie on any axis (3, 11), because both lines pass through this point (4, 14), because one of the lines passes through this point (4, 14), because the point lies between the two axes

OpenStudy (anonymous):

@Mehek14 @babykel

OpenStudy (anonymous):

honestly this whole lessons getting me stressed you explain i get it but then i go on to the next and im like errr

OpenStudy (anonymous):

@jim_thompson5910

Mehek (mehek14):

sub line C in Line D for y \(x+8 = 3x+2\) combine like terms 3x - x = 2x 8 - 2 = 6 \(6=2x\) divide 6/2 = 3 \(3=x\)

Mehek (mehek14):

solution = when lines intersect so it goes through both of them

OpenStudy (anonymous):

so B

Mehek (mehek14):

yup

OpenStudy (anonymous):

can i post three all at once?

OpenStudy (anonymous):

all the same kind of stuff

Mehek (mehek14):

yea

OpenStudy (anonymous):

do you like math that much like wow i would have been like no ive helped enough haha

OpenStudy (anonymous):

your really smart

OpenStudy (anonymous):

Well thank lol jk

Mehek (mehek14):

yes I ♥ math :)

Mehek (mehek14):

it's pretty easy for me

OpenStudy (anonymous):

i do not see how @Mehek14

Mehek (mehek14):

?

OpenStudy (anonymous):

how dont see how math is easy @Mehek14

Mehek (mehek14):

idk i can easily understand math

OpenStudy (anonymous):

fr tho like woah haha @babykel sorryyy